0 0
Add a comment Improve this question Transcribed image text
Answer #1


Total Tenere are 14 instances و outlook | Rain Sunny I overcast outlook Temp= cool | Tempahot Tem=mild Sunny مران از - ا over outlook | Humidity=high H Humidity = normal alo llo sunny overcast Rain hole slupla da solo Temperature, Humidity play=yes IPLYES) - 5 9 Tú P (Nol= pf windweak tyle having instances inces Humidity = Hight num of Humidity = normal =) instances havingP ( High / No) = P(Nol High) P (High) P (NO) Pirosmal lygs) - PC Esimorvial] * p loov mal) Plyes) p(normal) P(Namal (No) = P(num of instances having Tema Hot =4 Plres/ Hot ) a num of instances having Tema Milda G P(yes/ Mild) = 4 mium of instances ha미 1001yes) =p[yes \cool) x PC cool) P4 회에티지 M/outlook - 3 ok = sunny=5 P(Ves / sunny num of instances having a - overcast-4 P(yes lovely - in of instances having outlook o* P(Rain) ܠܐ ܘܐܠܘܠܣܪ et p( Ranival yes) = P( res a 9 x foutlook-sunny temperature = shot humidity srovnal p( xl yes) = f(kes)! PL xl yes) = m(Yes) < p sunettes) xp hot (yes) 21x9 P(ANO) = P(NO) to some hos) * pfhot (No) te p (normal wo) ook 9.29 so t